Consider the following equations: SUPPLY: Q=10+2P DEMAND: Q=60-3P

b) If the government imposes a ceiling price (P max) of $8.50 per unit. Would it result in a shortage or a surplus or a surplus? Show in the graph.
Answer #1

Ans: Shortage

Explanation:

Quantity demand = 60 - 3P

Quantity supply = 10 + 2P

At equilibrium , Quantity demand = Quantity supply

60 - 3P = 10 + 2P

5P = 60 -10

P = 50 / 5 = 10

Equilibrium quantity = 10 + 2P = 10 + ( 2 *10) = 10 + 20 = 30

If the government imposes a ceiling price (P max) of $8.50 per unit then there will be shortage. It is shown in the following figure. At ceiling price (P max) of $8.50 per unit , quantity supply is less than quantity demand. It is because, the producers supply less at lower price ( less than equilibrium price ). As a result , there is shortage of that product.

1o QUANTITY (in UnidA